after plowing a border inside a rectangular field 60 rods wide and 80 rods long, a farmer finds that one half of the field remains to be plowed. how wide is the border?
Answer by jorel555(1290)

The field is 60 x 80, or 4800 square rods. The border is on all sides, so the area left to be plowed equals 60-2n by 80-2n. So:
(60-2n)(80-2n)=2400
4800-280n+4nē=2400
4nē-280n+2400=0
nē-70n+600=0
(n-60)(n-10)=0
n=60 or 10
This leaves us with a border 10 rods wide. ☺☺☺☺